### v1

@@ 2.1 $\ell$-isolated cliques
We remark that the property of $\ell$-isolation is monotone. Namely, if $\ell_{1}>\ell_{2}>0$, then any $\ell_{2}$-isolated set is also $\ell_{1}$-isolated. This follows immediately from the definition of $\ell$-isolation, since for an $\ell_{2}$-isolated set $I$ we have $\delta_{e}(I)<\ell_{2}\cdot k<\ell_{1}\cdot k$ when $\ell_{1}>\ell_{2}$. Hence, when searching for $\ell$-isolated maximal cliques, the larger is the value of $\ell$, the weaker is the isolation of the cliques relative to the rest of the graph, and the is larger the number of $\ell$-isolated cliques in the graph. If $\ell$ is high enough, the set of $\ell$-isolated maximal cliques coincides with the set of all maximal cliques of $G$.
@@ 2.2 Maximal clique enumeration with pivoting
6:  if $P=\varnothing$ and $X=\varnothing$ then return $C$
7:  Choose a pivot vertex $v_{p}$ in $P\cup X$ with highest $|N(v_{p})\cap P|$
8:  for all $v\in P\setminus N(v_{p})$ do
@@
10:   $P\leftarrow P\setminus\{v\}$
11:   $X\leftarrow X\cup\{v\}$
@@ 3 Pruning for $\ell$-isolated maximal clique enumeration
Pruning by upper bounding the maximum clique size. Lemma˜3.3 implies that a check on $\omega(G[P])$ would be sufficient to prune at node $r$. Unfortunately, computing $\omega(G[P])$ is difficult — max-clique being $\mathsf{NP}$-hard. Hence, following an approach suggested in [DBLP:conf/kdd/WangCF13] for the computation of $\tau$-visible summaries, we use upper bounds to $\omega(G[P])$ to guide pruning. We will prove in Theorem˜3.5 that this maintains correctness with respect to our $\ell$-isolation problem. In particular, $\omega(G[P])$ cannot be larger than:
@@ 4 Experimental setup
Performance metrics. In addition to running times, we profiled the number of recursive calls and contrasted these values with the number of $\ell$-isolated cliques. The running time of each heuristic depends indeed both on its pruning capabilities and on the overhead introduced by the pruning test in Equation˜4: the more complex the pruning strategy, the larger the test overhead.
@@ 5.1 On the distribution of $\ell$-isolated maximal cliques
In Figure˜1 we summarize the distribution of maximal cliques by size, together with their breakdown into $\ell$-isolated cliques for different values of $\ell$.
@@
This analysis can be refined by considering the color segments within each bar of Figure˜1, that show how many maximal cliques are $\ell$-isolated for several values of $\ell$. For small values of $\ell$ (e.g., $\ell\leq 20$), the isolation requirement is very strict and only a small fraction of small cliques qualify: large cliques are typically excluded, as they tend to be embedded in denser regions of the graph with many external connections. The distributions for small $\ell$ are thus sharply truncated and dominated by cliques with just a few vertices. As $\ell$ increases, the number of $\ell$-isolated cliques grows, being the property monotone (see Section˜2.1). In addition, larger values of $\ell$ allow increasingly larger cliques to appear in the summary, extending the tail of the distribution. For large thresholds, such as $\ell=200$, many more cliques are admitted: in many networks (e.g., brightkite, web-google, cit-hepph) the profile for 200-isolated cliques approaches that of the total distribution, making the isolation criterion almost ineffective.
@@ 5.2 Assessment of different pruning strategies
We now assess the effectiveness of different pruning strategies considering two different metrics: number of recursive calls and running time. Because the results are largely consistent across different values of $\ell$, Figure˜2 shows only the case $\ell=50$. The left radar plot in Figure˜2 reports the percentage of recursive calls with respect to the number one would have without pruning. This gives a quantitative view of the pruning power of the heuristics: the larger the area, the weaker the pruning. In spite of a large variability across benchmarks, size is always the weakest, as expected from the theoretical analysis, producing the largest number of calls. On some benchmarks, e.g., twitter, this number can be much larger than the other strategies and very close to the baseline. As expected, combo and softcore overlap perfectly, since combo is size followed by softcore whenever size fails, thus inheriting the pruning power of softcore. While degeneracy consistently achieves the strongest pruning, yielding the smallest areas across all datasets, it is rather similar to softcore and combo.
@@ 5.3 Comparison with state-of-the-art competitors
In Figure˜3 we compare our best heuristics, size and combo, against HKMN and TTT, also analyzing how $\ell$ affects the running times. We chose a range of values for $\ell$ based on the observations in Section˜5.1, covering both the lower and the upper spectrum of the isolation factor. A few trends emerge clearly from the charts. First, the running time of TTT is independent of $\ell$, thus remaining constant. The two heuristics size and combo typically deliver the best performance over a large range of values of $\ell$, with runtimes that remain low and scale smoothly as $\ell$ increases, when larger portions of $\mathcal{T}_{G}$ need to be explored. Their curves are often close (see, e.g., livemocha or cfinder-google), combo is marginally faster than size on a few instances (most notably, stanford and italian-cnr), but size seems preferable in most of the other ones. On the other side, HKMN is by far the slowest algorithm when $\ell$ gets large, with runtimes that grow quickly with $\ell$: this is due to an $\ell$-parameterized subroutine for finding minimal vertex covers. It can be instead competitive when $\ell$ is very small. TTT is only occasionally competitive with size and combo, but still outperforms HKMN for large $\ell$ except on web-google and dblp. The observed relative differences between the algorithms persist across the datasets, showing that even weak pruning heuristics that add a small test overhead on nodes of $\mathcal{T}_{G}$ translate into significant speedups that can be as large as $4\times$ with respect to HKMN and $2\times$ with respect to TTT.
@@ 5.4 Experiments with synthetic networks
We first discuss experiments on BAn,m networks. Figure˜4 compares TTT against HKMN (Figure˜4(a)) and size and combo (Figure˜4(b)), respectively. For $m=25$ and $\ell\leq 20$, we have that HKMN is comparable with combo and both of them perform better than size. However, the dashed red line also shows that in the above range the $\ell$-isolated maximal cliques are very few (we also experimentally assessed their small size, akin to real-world networks). For higher values of $\ell$, the running time of HKMN seems to be strongly dependent on the output size. Heuristic size (and sometimes combo), instead, provide a more efficient solution when $\ell\geq 30$. Curve trends are similar when $m=50$. The percentage of $\ell$-isolated cliques decreases: for $\ell=250$ it is barely above $60\%$ and remains close to zero for $\ell\leq 50$ (in this range HKMN is the fastest algorithm). When $\ell>50$, TTT is faster than combo and HKMN, but size remains the preferred solution.
Networks generated according to the $G_{n,m,p}$ model have substantially different properties: they are typically very dense and $\ell$-isolated maximal cliques are barely present for small values of $\ell$ (e.g., $\ell<80$). Conversely, for high values of $\ell$ (e.g., $\ell>100$), their percentage becomes close to $100\%$, after a very steep transition (see the dashed red line in Figure˜5). We reproduced the analysis of [DBLP:journals/tcs/HuffnerKMN09], reporting in Figure˜5 only the results with respect to $\ell$. Our results confirm those in [DBLP:journals/tcs/HuffnerKMN09], with HKMN outperforming other approaches for small values of $\ell$. TTT seems instead the algorithm of choice for larger values. In any case, these graphs appear to represent worst-case scenarios for the algorithms proposed in this paper, although they remain very far from the properties of real-world networks.

### v2

@@ 2.1 $\ell$-isolated cliques
We remark that the property of $\ell$-isolation is monotone. Namely, if $\ell_{1}>\ell_{2}>0$, then any $\ell_{2}$-isolated set is also $\ell_{1}$-isolated. This follows immediately from the definition of $\ell$-isolation, since for an $\ell_{2}$-isolated set $I$ we have $\delta_{e}(I)<\ell_{2}\cdot k<\ell_{1}\cdot k$ when $\ell_{1}>\ell_{2}$. Hence, when searching for $\ell$-isolated maximal cliques, the larger is the value of $\ell$, the weaker is the isolation of the cliques relative to the rest of the graph, and the larger is the number of $\ell$-isolated cliques in the graph. If $\ell$ is high enough, the set of $\ell$-isolated maximal cliques coincides with the set of all maximal cliques of $G$.
@@ 2.2 Maximal clique enumeration with pivoting
6:  if $P=\varnothing$ and $X=\varnothing$ then
7:   yield $C$
8:   return
9:  Choose a pivot vertex $v_{p}$ in $P\cup X$ with highest $|N(v_{p})\cap P|$
10:  for all $v\in P\setminus N(v_{p})$ do
@@
12:   $P\leftarrow P\setminus\{v\}$
13:   $X\leftarrow X\cup\{v\}$
@@
Figure 1: A running example of Algorithm˜1 on a small graph. Green, blue, and red vertices belong to $P$, $C$, and $X$, respectively. The current pivot vertex $v_{p}$ is circled with a dotted curve. The neighbourhood of $v_{p}$ is circled with a dashed curve, the current vertex $v$ is circled with a red solid curve.
A running example of Algorithm˜1 on a small graph is depicted in Figure˜1. When the first call of the procedure ProcMCE is launched, all the vertices are in $P$, while $X=C=\varnothing$ (Figure˜1(b)). The pivot vertex is $v_{p}=v_{2}$ (line 9 of Algorithm˜1) and the procedure iterates on vertex $v_{2}$ only, since all other vertices in $P$ are neighbors of $v_{p}$ (line 10 of Algorithm˜1). Hence, in the second call of ProcMCE (Figure˜1(c)) vertex $v_{2}$ has been moved to $C$, $v_{p}=v_{3}$, and line 10 of Algorithm˜1 only iterates on $v_{3}$. Instead, Call 3 iterates on both $v_{1}$ and $v_{4}$: the iteration on $v_{1}$ (Figure˜1(d)) launches Call 4 (Figure˜1(e)) while the iteration on $v_{4}$ (Figure˜1(f)) launches Call 5 (Figure˜1(g)). Call 4 and Call 5 detect each a maximal clique, since both $P$ and $X$ are empty (line 6 of Algorithm˜1). When Call 5 terminates, the control returns to Call 3 (Figure˜1(h)) and then to Call 2 (Figure˜1(i)) and Call 1 (Figure˜1(j)), terminating the computation.
@@ 3 Pruning for $\ell$-isolated maximal clique enumeration
Pruning by upper bounding the maximum clique size. Lemma˜3.3 implies that a check on $\omega(G[P])$ would be sufficient to prune at node $r$. Unfortunately, computing $\omega(G[P])$ is difficult — max-clique being $\mathsf{NP}$-hard. Hence, we exploit upper bounds to $\omega(G[P])$ to guide pruning, inspired by the approach suggested in [DBLP:conf/kdd/WangCF13] for the computation of $\tau$-visible summaries. We will prove in Theorem˜3.5 that this maintains correctness with respect to our $\ell$-isolation problem. In particular, $\omega(G[P])$ cannot be larger than:
@@ 4 Experimental setup
Performance metrics. In addition to running times, we profiled the number of recursive calls and contrasted these values with the number of $\ell$-isolated cliques. The running time of each heuristic depends indeed both on its pruning capabilities and on the overhead introduced by the pruning test in Equation˜4: the more complex the pruning strategy is, the larger the test overhead.
@@ 5.1 On the distribution of $\ell$-isolated maximal cliques
In Figure˜2 we summarize the distribution of maximal cliques by size, together with their breakdown into $\ell$-isolated cliques for different values of $\ell$.
@@
This analysis can be refined by considering the color segments within each bar of Figure˜2, that show how many maximal cliques are $\ell$-isolated for several values of $\ell$. For small values of $\ell$ (e.g., $\ell\leq 20$), the isolation requirement is very strict and only a small fraction of small cliques qualify: large cliques are typically excluded, as they tend to be embedded in denser regions of the graph with many external connections. The distributions for small $\ell$ are thus sharply truncated and dominated by cliques with just a few vertices. As $\ell$ increases, the number of $\ell$-isolated cliques grows, being the property monotone (see Section˜2.1). In addition, larger values of $\ell$ allow increasingly larger cliques to appear in the summary, extending the tail of the distribution. For large thresholds, such as $\ell=200$, many more cliques are admitted: in many networks (e.g., brightkite, web-google, cit-hepph) the profile for 200-isolated cliques approaches that of the total distribution, making the isolation criterion almost ineffective.
@@ 5.2 Assessment of different pruning strategies
We now assess the effectiveness of different pruning strategies considering two different metrics: number of recursive calls and running time. Because the results are largely consistent across different values of $\ell$, Figure˜3 shows only the case $\ell=50$. The left radar plot in Figure˜3 reports the percentage of recursive calls with respect to the number one would have without pruning. This gives a quantitative view of the pruning power of the heuristics: the larger the area, the weaker the pruning. In spite of a large variability across benchmarks, size is always the weakest, as expected from the theoretical analysis, producing the largest number of calls. On some benchmarks, e.g., twitter, this number can be much larger than the other strategies and very close to the baseline. As expected, combo and softcore overlap perfectly, since combo is size followed by softcore whenever size fails, thus inheriting the pruning power of softcore. While degeneracy consistently achieves the strongest pruning, yielding the smallest areas across all datasets, it is rather similar to softcore and combo.
@@ 5.3 Comparison with state-of-the-art competitors
In Figure˜4 we compare our best heuristics, size and combo, against HKMN and TTT, also analyzing how $\ell$ affects the running times. We chose a range of values for $\ell$ based on the observations in Section˜5.1, covering both the lower and the upper spectrum of the isolation factor. A few trends emerge clearly from the charts. First, the running time of TTT is independent of $\ell$, thus remaining constant. The two heuristics size and combo typically deliver the best performance over a large range of values of $\ell$, with runtimes that remain low and scale smoothly as $\ell$ increases, when larger portions of $\mathcal{T}_{G}$ need to be explored. Their curves are often close (see, e.g., livemocha or cfinder-google), combo is marginally faster than size on a few instances (most notably, stanford and italian-cnr), but size seems preferable in most of the other ones. On the other side, HKMN is by far the slowest algorithm when $\ell$ gets large, with runtimes that grow quickly with $\ell$: this is due to an $\ell$-parameterized subroutine for finding minimal vertex covers. It can be instead competitive when $\ell$ is very small. TTT is only occasionally competitive with size and combo, but still outperforms HKMN for large $\ell$ except on web-google and dblp. The observed relative differences between the algorithms persist across the datasets. Even weak pruning heuristics that add a small test overhead on nodes of $\mathcal{T}_{G}$ translate into significant speedups, which can be as large as $4\times$ and $2\times$ with respect to HKMN and TTT, respectively.
@@ 5.4 Experiments with synthetic networks
We first discuss experiments on BAn,m networks. Figure˜5 compares TTT against HKMN (Figure˜5(a)) and size and combo (Figure˜5(b)), respectively. For $m=25$ and $\ell\leq 20$, we have that HKMN is comparable with combo and both of them perform better than size. However, the dashed red line also shows that in the above range the $\ell$-isolated maximal cliques are very few (we also experimentally assessed their small size, akin to real-world networks). For higher values of $\ell$, the running time of HKMN seems to be strongly dependent on the output size. Heuristic size (and sometimes combo), instead, provide a more efficient solution when $\ell\geq 30$. Curve trends are similar when $m=50$. The percentage of $\ell$-isolated cliques decreases: for $\ell=250$ it is barely above $60\%$ and remains close to zero for $\ell\leq 50$ (in this range HKMN is the fastest algorithm). When $\ell>50$, TTT is faster than combo and HKMN, but size remains the preferred solution.
Networks generated according to the $G_{n,m,p}$ model have substantially different properties: they are typically very dense and $\ell$-isolated maximal cliques are barely present for small values of $\ell$ (e.g., $\ell<80$). Conversely, for high values of $\ell$ (e.g., $\ell>100$), their percentage becomes close to $100\%$, after a very steep transition (see the dashed red line in Figure˜6). We reproduced the analysis of [DBLP:journals/tcs/HuffnerKMN09], reporting in Figure˜6 only the results with respect to $\ell$. Our results confirm those in [DBLP:journals/tcs/HuffnerKMN09], with HKMN outperforming other approaches for small values of $\ell$. TTT seems instead the algorithm of choice for larger values. In any case, these graphs appear to represent worst-case scenarios for the algorithms proposed in this paper, although they remain very far from the properties of real-world networks.